SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Feb.28, 2006)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su

Operations

  • The vacuum activity near the IP is still there after the replacement of the Q4 chamber last week. Selina is working on locating the vacuum leak with L1 accepts in colliding beams and background single beam runs. Selina showed some plots here. The structure at z=+2.3m is worth looking into since it also agrees with the average value of other analyses from Brian Petersen's talk for the collobration meeting last week. Su Dong and Rainer made some suggestions to study the events in that region.
  • One of the L3 nodes was trimmed on Friday evening during cosmics running. It was recovered in the next run.
  • There was high L1 trigger rate in cosmic runs last Friday night. This is partly due to the the trigger mask and the "avalanche" mode IFR was running. Changing the trigger mask to use EMT triggers avoided the high rate.
  • Last Saturday, we got paged because the GLT objects 8-16 had zero entries, but they were expected due to the cosmic_nodch configuration. Rahmat modified the JAS page to include more information on the GLT objects so that the shifters can identify which objects are expected to be present for the different running configurations.
  • The long-term instruction for pilots was updated with the TSF-X power cycle information.

DCT Online

  • Fang is working on the L1Accept codes to ship out input datagram when there are FEX errors.

L3

  • Rainer mentioned that we should deploy the version of the L3 tracking code that removes the tracking bias on the high-occupancy events (see HN posted). Another motivation for putting in a new version now is to add a new z0 position plot which extends the monitoring into the vacuum problem area. He will talk to the online folks about this.

Data Quality and Monitoring

  • Rahmat showed some plots for the stripcharts for all events (hadronics + mu-pairs). There seems to be some shifts in the z0 residual. This looks like related to a shift in the beam spot. See Su Dong's HN posting after the meeting.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.